# Bulk-Boundary Correspondence and Exceptional Points for a Dimerized Hatano-Nelson Model with Staggered Potentials

### v1

@@ VI Entanglement entropy
In this section we investigate the entanglement entropy for the non-Hermitian model (II) with periodic boundary conditions. We use the reduced density matrix approach for free fermions [52, 53] which has been extended to the non-Hermitian realm [27, 28, 54, 26, 55] using the generalized density matrix $\rho=|r\rangle\langle\ell|$ where $\langle\ell|$ and $|r\rangle$ are bi-orthogonal left and right eigenstates of (II). The key quantity is the two-point correlation matrix with elements
$C_{ij}=\langle\ell|c_{i}^{\dagger}c_{j}|r\rangle={\rm tr}\,\left(\rho c_{i}^{\dagger}c_{j}\right)$ (VI.1)
@@
where $\nu_{j}$ are the eigenvalues of the correlation matrix (VI.1) restricted to $\mathcal{A}$. In this paper we only consider $\mathcal{A}=\{1,\dots,N/2\}$, that is, we cut the system in half. For the model considered in this paper, in contrast to Hermitian systems, the eigenvalues $\nu_{j}$ and $S_{\mathcal{A}}$ can be complex. Nevertheless, we will indiscriminately call $S_{\mathcal{A}}$ the entanglement entropy. In addition, we choose the branch cut of the logarithm to be on the negative real axis.
@@
We now investigate the spectrum of the correlation matrix (VI.12) and the associated entanglement entropy for indices in (VI.12) restricted to $i,j\in\{1,\dots,N/2\}$. We perform numerical calculations for large lattices, in which case we only keep eigenvalues of the correlation matrix that satisfy $|\nu_{j}|>10^{-50}$ and $|1-\nu_{j}|>10^{-50}$. Additionally, we set to zero real or imaginary parts of $\nu_{j}$ and $1-\nu_{j}$ that are individually smaller than $10^{-50}$. This is particular important in cases where the eigenvalues of the correlation matrix are negative with a tiny imaginary part, whose sign can numerically oscillate around the branch cut. Before considering numerics, as a warm-up, we study the trivial 2-cell model ($N=4$) analytically. Both zero field $u=0$ and $\mathcal{PT}$-symmetric cases are considered.
@@ VI.1 Sub-lattice symmetric case
For the zero field case, setting $V_{R}=W_{R}=1$, let us consider transitions between the phases $(\nu_{1},\nu_{2})=(1,-1)$ (with $W_{L}>1$) and $(0,-1)$, between $(1,0)$ and $(0,0)$ and between $(1,-1)$ (with $W_{L}<-1$) and $(0,-1)$, recall the phase diagram shown in the left panel of Fig. 2. For concreteness, let us consider only $V_{L}\geq 0$. Recall that the spectrum of $\mathcal{T}_{N}$ in the phases $(1,-1)$ and $(0,0)$ is in general composed of two lobes in the complex plane either separated by the imaginary axis or real axis, therefore being characterized by $I_{1}=0$. On the other hand, the spectrum in the phases $(1,0)$ and $(0,-1)$ traces a closed path around $E_{B}=0$, therefore being characterized by $I_{1}\neq 0$.
We start with the 2-cell model ($N=4$). For this lattice size, there is no lobe in the spectrum, and the eigenvalues are either purely real or purely imaginary. We fill the states with energy $-\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$ and leave empty the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$. The associated correlation matrix is given by,
@@
We observe that within the phases characterized by $(1,-1)$ the eigenvalues form conjugate pairs of the form $1/2\pm\text{i}\alpha$, see panels (a) and (c) in Fig. 9 for the fixed values $W_{L}=\pm 1.5$. The same form of eigenvalues is observed for other values of $|W_{L}|>1$. Thus, in this phase, the entanglement entropy is real and given by,
@@
Increasing $V_{L}$ and crossing the transition point $V_{L}=1$, we observe that the eigenvalues of the correlation matrix in the phase $(0,-1)$ are complex, with the form $\beta_{1,2}\pm\text{i}\alpha$. That is, the eigenvalues have imaginary parts with opposite signs but with different real parts $\beta_{2}=1-\beta_{1}$. This leads to complex values for the entanglement entropy. However, for large $|W_{L}|$ the imaginary part tends to disappear when $V_{L}$ increases. In the limit $V_{L}\rightarrow\infty$ and $|W_{L}|\rightarrow\infty$, the exact eigenvalues of the correlation matrix are $41/2\pm\sqrt{2}/$ 4, producing a real entropy $S_{2}\approx 0.832991$.
@@
In the phase $(1,0)$, the eigenvalues once more have the form $\beta_{1,2}\pm\text{i}\alpha$ with $\beta_{2}=1-\beta_{1}$, and the entanglement entropy is complex (except if $V_{L}=W_{L}$), see panel (b) in Fig. 9 for the fixed value $W_{L}=0.1$. Moving to the phase $(0,0)$, we observe that the eigenvalues of $C$ are real, however, they are greater than unity or smaller then zero, and therefore also produce a complex entanglement entropy. Only in the limit $V_{L}\rightarrow\infty$ and $W_{L}\rightarrow 0$, do the eigenvalues of the correlation matrix tend to $1$ and 00, thus producing a zero entanglement entropy.
@@
Figure 9: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the trivial two-cell model ($N=4$) with zero field $u=0$, $W_{R}=V_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)-(c) as a function of $V_{L}$.
@@
Now, we consider larger lattice sizes using numerical diagonalization. We observe a variety of possible types of eigenvalues in the correlation spectrum, similar to the trivial 2-cell model. In some phases, we may find eigenvalues with $\text{Re}(\nu_{j})>1$ as well as $\text{Im}(\nu_{j})\neq 0$. Also, we note that near the transition points the behaviour of the correlation matrix eigenvalues is more intricate. Let us analyze each phase in details. Results are collected in Fig. 10 for a lattice composed of $N=120$ sites. In the insets of Fig. 10, we show the typical form of the complex spectrum indicating the filling adopted, as well as the real part of the eigenvalues of the correlation matrix, considering a smaller lattice with $N=52$ for better visualization of the results.
@@
Figure 10: Entanglement entropy for the model with sublattice symmetry (zero field $u=0$) for $V_{R}=W_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)–(c), as a function of $V_{L}$. The lattice size is $N=120$. In the insets, the typical form of the complex spectrum is shown; the occupied states are indicated by filled circles while the unoccupied states by open circles. The typical eigenvalues of the correlation matrix are represented by empty up triangles (only the real part is shown since the complex part when present is small). For better visualization, in the insets we consider $N=52$.
[…]
@@
Moving to the phases $(0,-1)$, see once more panels (a) and (c) in Fig. 10, we note that all eigenvalues of the correlation matrix satisfy $10<\text{Re}(\nu_{j})<$ 1, with most eigenvalues being either close to unity or zero. Some intermediate eigenvalues appear, and these carry a small imaginary part. They produce a finite entanglement entropy, also complex, but with a small imaginary part. However, the number of intermediate states actually grows with the length of the chain. This implies a scaling of the entanglement entropy, see Fig. 11, where we fix $W_{L}=1.5$ and some values of $V_{L}$ deep inside the phase. We observe a scaling of the form $\text{Re}(S_{N})\sim(c/3)\log(N)$ with $c=1$, as well as a decay $\text{Im}(S_{N})\sim\exp(-\log(N))$ of the imaginary part. It follows that, remarkably, this phase may be described by a CFT with central charge $c=1$, even though it has a point gapped spectrum. Similar results can be obtained for $W_{L}<-1$ in the same phase $(0,-1)$.
@@
Figure 11: Scaling of the entanglement entropy in the phase $(0,-1)$ for $W_{L}=1.5$ and values of $V_{L}$ deep inside the phase.
@@
The phase $(1,0)$, see panel (b) in Fig. 10, is similar to the phases $(0,-1)$. The scaling of the entanglement entropy for this case is shown in Fig. 12. Also in this phase the scaling of the entanglement entropy is consistent with a central charge $c=1$.
@@ VI.2 $\mathcal{PT}$-symmetric case
Let us now consider the $\mathcal{PT}$ symmetric case with PBC, setting $u=1$. For simplicity, we will restrict ourselves to the quadrant with $V\geq 0$ and $W\geq 0$ in the phase diagram shown in the right panel of Fig. 2. Then, four regions must be analyzed. The region $|W-V|<1<W+V$ is the exceptional phase, characterized by a large number of exceptional points (recall Sec. III.2). In this phase, the $\mathcal{PT}$-symmetry is broken. The regions $W-V<-1$ and $W-V>1$ are adiabatically connected to the standard SSH model via $u\rightarrow 0$. Then, the former is called the $\mathcal{PT}$-unbroken trivial phase and the latter the $\mathcal{PT}$-unbroken topological phase. In the phase $W+V<1$ the spectrum is purely imaginary, and we name it the complex phase.
As before, we start our analysis by the simple 2-cell model ($N=4$), which has the following correlation matrix,
@@
associated with the filled states $-\sqrt{(V\mp W)^{2}-1}$ and empty states $\sqrt{(V\mp W)^{2}-1}$. We plot in Fig. 13 the eigenvalues of $C$ and the associated entanglement entropy for some fixed values of $V$ as a function of $W$, covering all the four regions of the phase diagram.
For $V=0$ and $10<W<$ 1 that is, within the complex phase, see panel (a) in Fig. 13, we can observe that the eigenvalues of the correlation matrix are real with $\nu_{j}>1$ or $\nu_{j}<0$ until the phase transition point $W=1$ is reached. As a consequence, the entanglement entropy is complex in this phase. Crossing the transition point ($W=1$) towards the $\mathcal{PT}$-unbroken topological phase, we note that the eigenvalues have the form $1/2\pm\text{i}\alpha$, and the entanglement entropy is given by (VI.14). In the limit $W\rightarrow\infty$, $\alpha$ becomes negligible and the entanglement entropy tends to $S_{2}\rightarrow 2\log(2)$, indeed characteristic of a topological phase with winding $I=1$.
Increasing $V$ to $V=0.5$, see panel (b) in Fig. 13, the behaviour of the eigenvalues $\nu_{j}$ is the same as for $V=0$ until the exceptional phase is reached at $W=0.5$. Within the exceptional phase, the eigenvalues $\nu_{j}$ have the form $\beta_{1,2}\pm\text{i}\alpha$, that is, the eigenvalues have the same imaginary part with opposite signs but different real parts related by $\beta_{2}=1-\beta_{1}$. As a consequence, the entanglement entropy is complex in the exceptional phase. This phase ends at $W=1.5$ from which point the $\mathcal{PT}$-unbroken topological phase is again reached being characterized by a real entanglement entropy which tends to $S_{2}\rightarrow 2\log(2)$ deep inside the phase.
To conclude the analysis of the 2-cell model, we consider $V=1.5$ such that the $\mathcal{PT}$-unbroken trivial phase is considered. This phase for $V=1.5$ occurs when $0.50<W<$ 0.5, and we note that there the eigenvalues are real with $\nu_{j}>1$ or $\nu_{j}<0$, producing again a complex entanglement entropy except if $W=0$. However, deep inside the phase ($V\rightarrow\infty$), the eigenvalues tend to 00 and $1$, leading to a null entanglement, as expected from a trivial phase. For $2.50.5<W<$ 2.5 we encounter once more the exceptional phase and for $W>2.5$ the $\mathcal{PT}$-unbroken topological phase.
Figure 13: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the trivial $\mathcal{PT}$-symmetric two-cell model ($N=4$) with $u=1$, some fixed values of $V$ indicated in the labels (a)-(c) as a function of $W$.
@@
First, let us consider $V=0$ and vary $W$. For $10<W<$ 1, we note that the correlation matrix has real eigenvalues, most of them 00 or $1$, but there is one eigenvalue $\nu_{\text{max}}>1$ and one eigenvalue $\nu_{\text{min}}<0$. Remarkably, these eigenvalues do not depend on the lattice size, and therefore correspond to the two eigenvalues of the 2-cell model. For instance, for $V=0,W=0.5$, the eigenvalues are $\nu_{\text{max}}\approx 1.077350$ and $\nu_{\text{min}}\approx-0.077350$, which produces a finite entanglement entropy with negative real part and positive imaginary part, precisely the same as the 2-cell model, recall panel (a) in Fig. 13. Crossing the transition point $W=1$, we note that most of the eigenvalues of the correlation matrix are 00 or $1$, but there are a couple of eigenvalues of the form $1/2\pm\text{i}\alpha$, where $\alpha$ also does not depend on the lattice size. However, we note that for a fixed length $\alpha\rightarrow 0$ when $W\rightarrow\infty$. Therefore, like in some of the previously considered cases, the entanglement approaches $S_{N/2}\rightarrow 2\log(2)$ deep inside this phase.
@@
Now we consider $V\neq 0$. Similarly to the 2-cell model, by fixing $10<V<$ 1, we can reach the complex phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase by varying $W$. On the other hand, by fixing $V>1$ and varying $W$ we can reach the $\mathcal{PT}$-unbroken trivial phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase.
@@
Let us consider once more $V=0.5$. Within the complex phase ($0.50<W<$ 0.5 for $V=0.5$), all the eigenvalues of the correlation matrix jump above unity or below zero, see panel (a) in Fig. 14 where the parameters $W=0.45$ and $V=0.5$ are used. This results in a complex value for $S_{N/2}$, see panel (a) in Fig. 15. We note that two eigenvalues stand out, being significantly greater than $1$ or less than 00. All the eigenvalues of the correlation matrix in this phase seem to vary slowly with the size of the chain, actually tending to stabilize. For example, the difference between the largest eigenvalues for the lattices with $N=52$ and with $N=120$ is of the order of $10^{-9}$. Therefore, the entanglement entropy in this phase seems to be finite. However, it is difficult to verify this property since the model parameters in this phase lie in a finite interval. Also, the phase is surrounded by critical lines that drive the entanglement entropy to high values.
@@
Moving to the $\mathcal{PT}$-unbroken topological phase, we note that the correlation spectrum is composed by eigenvalues close to $1$ and 00, with some of them being slightly bigger than $1$ or smaller then 00, as well as a conjugate pair of the form $1/2\pm\text{i}\alpha$, see panel (b) in Fig. 14 where the parameters are set to $W=1.7,V=0.5$. As the lattice size increases, we note that the eigenvalues tend to $1$ or 00, while $\alpha$ decreases slowly and seems to stabilize with the length of the chain. The associated entanglement entropy is real and finite. Going deep into the phase by increasing $W$, as before, $\alpha$ decreases and the entanglement tends to $S_{N/2}\rightarrow 2\log{2}$, see both panels (a) and (b) in Fig. 15.
@@
In the $\mathcal{PT}$-unbroken trivial phase, the correlation spectrum is also composed mostly of 00’s and $1$’s, some of them slightly bigger than $1$ or smaller than 00. Also, some eigenvalues carry an imaginary part in the form of conjugate pairs, see panel (d) in Fig. 14 for the model parameters $W=0.5$ and $V=1.8$. Interestingly, similarly to previous cases, both real and imaginary part of the eigenvalues vary slowly with the length of the chain. Nevertheless, deep inside this phase (large $V$), the eigenvalues tend to 00 or $1$ and produces zero entanglement, as expected from a trivial phase; see panel (b) in Fig. 15.
@@
Let us briefly discuss the exceptional phase. In this phase we observe an intricate behaviour of the correlation spectrum, characterized by various complex eigenvalues, see panel (c) in Fig. 14 where the parameters are set to $V=0.5,\,W=1.2$. This behaviour is reflected in an intricate evolution of the entanglement entropy, see panels (a) and (b) in Fig. 15. Interestingly, the entanglement entropy detects the presence of the real exceptional points (indicated by dashed lines in Fig. 15). For clarity, we only pick a few values of $W$ between the exceptional points. Note that in the thermodynamic limit the entanglement entropy is ill defined in the exceptional phase, since it is filled with exceptional points. For finite systems, we expect a scaling of $S_{N/2}$. This is numerically challenging and goes beyond the scope of this paper.

### v3

@@ VI Entanglement entropy
For a system in a topological phase, we expect that a non-zero lower bound for the entanglement entropy exists. This entanglement is protected and cannot be removed [60]. To study this aspect of topology, we investigate in this section the entanglement entropy for the non-Hermitian model (II) with periodic boundary conditions. The other novel aspect which motivates our study of the entanglement entropy is the $\mathcal{PT}$-symmetric phase in which exceptional points are dense. As we will show, this leads to an entanglement entropy which is nowhere continuous.
We use the reduced density matrix approach for free fermions [61, 62] which has been extended to the non-Hermitian realm [28, 29, 30, 31, 32, 33, 34, 35, 63] using the generalized density matrix $\rho=|r\rangle\langle\ell|$ where $\langle\ell|$ and $|r\rangle$ are bi-orthogonal left and right eigenstates of (II). The key quantity is the two-point correlation matrix with elements
$C_{ij}=\langle\ell|c_{i}^{\dagger}c_{j}|r\rangle={\rm tr}\,\left(\rho c_{i}^{\dagger}c_{j}\right)\,.$ (VI.1)
@@
where $\nu_{j}$ are the eigenvalues of the correlation matrix (VI.1) restricted to $\mathcal{A}$. In this paper we only consider $\mathcal{A}=\{1,\dots,N/2\}$, that is, we cut the system in half. For the model considered in this paper, in contrast to Hermitian systems, the eigenvalues $\nu_{j}$ and therefore also $S_{\mathcal{A}}$ can be complex. Nevertheless, we will indiscriminately call $S_{\mathcal{A}}$ the entanglement entropy. In addition, we choose the branch cut of the logarithm to be on the negative real axis. Before we proceed further, we might ask why $S_{\mathcal{A}}$ is complex and what the meaning of a complex entanglement entropy is. For the first part of the question it is important to remember that a non-Hermitian Hamiltonian is an effective approach obtained from a Master equation by dropping terms which induce quantum jumps. The density matrix in the proper Master equation is always positive semi-definite. $S_{\mathcal{A}}$ being complex is an artefact of the approximation. Nevertheless, the entanglement entropy can be a useful indicator for the correlations in the system and we expect, in particular, that it is sensitive to phase transitions and exceptional points. Physically, one can typically understand the real part of $S_{\mathcal{A}}$ as being still associated with the spreading of Schmidt values while the imaginary part is related to dissipation.
@@
We now investigate the spectrum of the correlation matrix (VI.12) and the associated entanglement entropy for indices in (VI.12) restricted to $i,j\in\{1,\dots,N/2\}$. We perform numerical calculations for large lattices, in which case we only keep eigenvalues of the correlation matrix that satisfy $|\nu_{j}|>\Lambda$ and $|1-\nu_{j}|>\Lambda$. Additionally, we set to zero real or imaginary parts of $\nu_{j}$ and $1-\nu_{j}$ that are individually smaller in magnitude than the chosen cutoff $\Lambda=10^{-50}$. This is particularly important in cases where the eigenvalues of the correlation matrix are negative with a tiny imaginary part, whose sign can numerically oscillate around the branch cut. By calculating the discarded weight $\sum_{|\nu_{j}|<\Lambda}|\nu_{j}|+\sum_{|1-\nu_{j}|<\Lambda}|1-\nu_{j}|$ we make sure that the cutoff does not affect the results on the scale shown in the figures provided that we stay on one branch of the logarithm. Before considering numerical results for large systems we will also study, as a warm-up, the 2-cell model ($N=4$) analytically. Both zero field $u=0$ and $\mathcal{PT}$-symmetric cases are considered.
@@ VI.1 Sub-lattice symmetric case
For the zero field case, setting $V_{R}=W_{R}=1$, let us consider transitions between the phases (i) $(\nu_{1},\nu_{2})=(1,-1)$ (with $W_{L}>1$) and (0,1)(0,-1), (ii) between $(1,0)$ and $(0,0)$, and (iii) between (1,1)(1,-1) (with $W_{L}<-1$) and (0,1)(0,-1), recall the phase diagram shown in the left panel of Fig. 2. For concreteness, let us consider only $V_{L}\geq 0$. Recall that the spectrum of $\mathcal{T}_{N}$ in the phases (1,1)(1,-1) and $(0,0)$ is in general composed of two lobes in the complex plane either separated by the imaginary axis or real axis, therefore being characterized by $I_{1}=0$. On the other hand, the spectrum in the phases $(1,0)$ and (0,1)(0,-1) traces a closed path around $E_{B}=0$, therefore being characterized by $I_{1}\neq 0$.
We start with the 2-cell model ($N=4$). For this lattice size, there is no lobe in the spectrum, and the eigenvalues are either purely real or purely imaginary. We fill the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$ and leave empty the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$. The associated correlation matrix is given by,
@@
We observe that within the phases characterized by (1,1)(1,-1) the eigenvalues form conjugate pairs of the form $1/2\pm\text{i}\alpha$, where $\alpha$ is a given function of $V_{L}$ and $W_{L}$, see panels (a) and (c) in Fig. 9 for the fixed values $W_{L}=\pm 1.5$. The same form of eigenvalues is observed for other values of $|W_{L}|>1$. Thus, in this phase, the entanglement entropy is real and given by,
@@
Increasing $V_{L}$ and crossing the transition point $V_{L}=1$, we observe that the eigenvalues of the correlation matrix in the phase (0,1)(0,-1) are complex, with the form $\beta_{1,2}\pm\text{i}\alpha$. That is, the eigenvalues have imaginary parts with opposite signs but with different real parts $\beta_{2}=1-\beta_{1}$. This leads to complex values for the entanglement entropy. However, for large $|W_{L}|$ the imaginary part tends to disappear when $V_{L}$ increases. In the limit $V_{L}\rightarrow\infty$ and $|W_{L}|\rightarrow\infty$, the exact eigenvalues of the correlation matrix are $1/2\pm\sqrt{2}/4$, producing a real entropy $S_{2}\approx 0.832991$.
@@
In the phase $(1,0)$, the eigenvalues once more have the form $\beta_{1,2}\pm\text{i}\alpha$ with $\beta_{2}=1-\beta_{1}$, and the entanglement entropy is complex (except if $V_{L}=W_{L}$), see panel (b) in Fig. 9 for the fixed value $W_{L}=0.1$. Moving to the phase $(0,0)$, we observe that the eigenvalues of $C$ are real, however, they are greater than unity or smaller than zero, and therefore also produce a complex entanglement entropy. Only in the limit $V_{L}\rightarrow\infty$ and $W_{L}\rightarrow 0$, do the eigenvalues of the correlation matrix tend to $1$ and 0, thus producing a zero entanglement entropy.
@@
Figure 9: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the two-cell model ($N=4$) with zero field $u=0$, $W_{R}=V_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)-(c) as a function of $V_{L}$. $S_{2}$ is discontinuous at the phase transitions.
@@
Now, we consider larger lattice sizes using numerical diagonalization. We observe a variety of possible types of eigenvalues in the correlation spectrum, similar to the 2-cell model. In some phases, we may find eigenvalues with $\text{Re}(\nu_{j})>1$ as well as $\text{Im}(\nu_{j})\neq 0$. Also, we note that near the transition points the behavior of the correlation matrix eigenvalues is more intricate. Let us now analyze each phase in detail. Results are collected in Fig. 10 for a lattice composed of $N=120$ sites. In the insets of Fig. 10, we show the typical form of the complex spectrum indicating the filling adopted, as well as the real part of the eigenvalues of the correlation matrix, considering a smaller lattice with $N=52$ for better visualization of the results.
@@
Figure 10: Entanglement entropy for the model with sublattice symmetry (zero field $u=0$) for $V_{R}=W_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)–(c), as a function of $V_{L}$. The lattice size is $N=120$. $S_{N/2}$ is discontinuous at the phase transitions. In the insets, the typical form of the complex spectrum is shown; the occupied states are indicated by filled circles while the unoccupied states by open circles. The typical eigenvalues of the correlation matrix are represented by empty up triangles (only the real part is shown since the complex part when present is small). For better visualization, in the insets we consider $N=52$.
Starting with the phases (1,1)(-1,1), see panels (a) and (c) in Fig. 10 for the fixed values $W_{L}=\pm 1.5$, we observe that most of the eigenvalues of the correlation matrix are either close to $1$ or 0. Some of the eigenvalues have the property $\nu_{j}>1$ or $\nu_{j}<0$, and there are two eigenvalues sitting at $1/2\pm\text{i}\alpha$. As the lattice size increases, we observe that those eigenvalues with $\nu_{j}>1$ or $\nu_{j}<0$ tend to unity or zero, while $\alpha$ decreases. For instance, for $N=52$, the lattice size considered in the insets, the highest eigenvalue is $\nu_{\text{max}}\approx 1.000000000290214$, the lowest $\nu_{\text{min}}=1-\nu_{\text{max}}$ and $\alpha\approx 4.275663\times 10^{-4}$, while for $N=120$ the highest eigenvalue for $W_{L}=1.5$ and $V_{L}=0.5$ is $\nu_{\text{max}}\approx 1.00000000000000000265940$, the lowest $\nu_{\text{min}}=1-\nu_{\text{max}}$ and $\alpha\approx 2.955440\times 10^{-7}$. Therefore, the entanglement entropy comes essentially from the conjugate eigenvalues $1/2\pm\text{i}\alpha$, and its value is also given by (VI.14). For a fixed length of the chain, but going deeper into the phase by increasing $|W_{L}|$, we also note that $\alpha$ decreases. Then, in the limit $\alpha\rightarrow 0$, the entanglement entropy has the finite value $S_{N/2}\rightarrow 2\log 2$ deep inside the phase. Similar considerations can be made for the phase (1,1)(1,-1) with $W_{L}<-1$; note however that a different filling must be used in this case.
@@
Moving to the phases (0,1)(0,-1), see once more panels (a) and (c) in Fig. 10, we note that all eigenvalues of the correlation matrix satisfy $0<\text{Re}(\nu_{j})<1$, with most eigenvalues being either close to unity or zero. Some intermediate eigenvalues appear, and these carry a small imaginary part. They produce a finite entanglement entropy, also complex, but with a small imaginary part. However, the number of intermediate states actually grows with the length of the chain. This implies a scaling of the entanglement entropy, see Fig. 11, where we fix $W_{L}=1.5$ and some values of $V_{L}$ deep inside the phase. We observe a scaling of the form $\text{Re}(S_{N})\sim(c/3)\log(N)$ with $c=1$, as well as a decay $\text{Im}(S_{N})\sim\exp(-\log(N))$ of the imaginary part. It follows that, remarkably, this phase may be described by a CFT with central charge $c=1$, even though it has a point gapped spectrum. Similar results can be obtained for $W_{L}<-1$ in the same phase (0,1)(0,-1).
@@
Figure 11: Scaling of the entanglement entropy in the phase (0,1)(0,-1) for $W_{L}=1.5$ and values of $V_{L}$ deep inside the phase.
@@
The phase $(1,0)$, see panel (b) in Fig. 10, is similar to the phases (0,1)(0,-1). The scaling of the entanglement entropy for this case is shown in Fig. 12. Also in this phase the scaling of the entanglement entropy is consistent with a central charge $c=1$.
@@ VI.2 $\mathcal{PT}$-symmetric case
Let us now consider the $\mathcal{PT}$ symmetric case with PBC, setting $u=1$. For simplicity, we will restrict ourselves to the quadrant with $V\geq 0$ and $W\geq 0$ in the phase diagram shown in the right panel of Fig. 2. Then, four regions must be analyzed.
As before, we start our analysis with the simple 2-cell model ($N=4$), which has the following correlation matrix,
@@
associated with the filled states $\sqrt{(V\mp W)^{2}-1}$ and empty states $\sqrt{(V\mp W)^{2}-1}$. We plot in Fig. 13 the eigenvalues of $C$ and the associated entanglement entropy for some fixed values of $V$ as a function of $W$, covering all the four regions of the phase diagram.
For $V=0$ and $0<W<1$ that is, within the complex phase, see panel (a) in Fig. 13, we observe that the eigenvalues of the correlation matrix are real with $\nu_{j}>1$ or $\nu_{j}<0$ until the phase transition point $W=1$ is reached. As a consequence, the entanglement entropy is complex in this phase. Crossing the transition point ($W=1$) towards the $\mathcal{PT}$-unbroken topological phase, we note that the eigenvalues have the form $1/2\pm\text{i}\alpha$, and the entanglement entropy is given by (VI.14). In the limit $W\rightarrow\infty$, $\alpha$ becomes negligible and the entanglement entropy tends to $S_{2}\rightarrow 2\log(2)$, indeed characteristic of a topological phase with winding $I=1$.
Increasing $V$ to $V=0.5$, see panel (b) in Fig. 13, the behavior of the eigenvalues $\nu_{j}$ is the same as for $V=0$ until the exceptional phase is reached at $W=0.5$. Within the exceptional phase, the eigenvalues $\nu_{j}$ have the form $\beta_{1,2}\pm\text{i}\alpha$, that is, the eigenvalues have the same imaginary part with opposite signs but different real parts related by $\beta_{2}=1-\beta_{1}$. As a consequence, the entanglement entropy is complex in the exceptional phase. This phase ends at $W=1.5$ from which point the $\mathcal{PT}$-unbroken topological phase is again reached being characterized by a real entanglement entropy which tends to $S_{2}\rightarrow 2\log(2)$ deep inside the phase.
To conclude the analysis of the 2-cell model, we consider $V=1.5$ such that the $\mathcal{PT}$-unbroken trivial phase is considered. This phase for $V=1.5$ occurs when $0<W<0.5$, and we note that there the eigenvalues are real with $\nu_{j}>1$ or $\nu_{j}<0$, producing again a complex entanglement entropy except if $W=0$. However, deep inside the phase ($V\rightarrow\infty$), the eigenvalues tend to 0 and $1$, leading to a vanishing entanglement entropy, as expected from a trivial phase. For $0.5<W<2.5$ we encounter once more the exceptional phase and for $W>2.5$ the $\mathcal{PT}$-unbroken topological phase.
Figure 13: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the $\mathcal{PT}$-symmetric two-cell model ($N=4$) with $u=1$, and fixed values of $V$ as indicated in the labels (a)-(c) as a function of $W$. $S_{2}$ is discontinuous at the phase transitions.
@@
Similarly to the 2-cell model, by fixing $0<V<1$, we can reach the complex phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase by varying $W$. On the other hand, by fixing $V>1$ and varying $W$ we can reach the $\mathcal{PT}$-unbroken trivial phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase.
@@
Let us consider once more $V=0.5$. Within the complex phase ($0<W<0.5$ for $V=0.5$), all the eigenvalues of the correlation matrix jump above unity or below zero, see panel (a) in Fig. 14 where the parameters $W=0.45$ and $V=0.5$ are used. This results in a complex value for $S_{N/2}$, see panel (a) in Fig. 15. We note that two eigenvalues stand out, being significantly greater than $1$ or less than 0. All the eigenvalues of the correlation matrix in this phase seem to vary slowly with the size of the chain, actually tending to stabilize. For example, the difference between the largest eigenvalues for the lattices with $N=52$ and with $N=120$ is of the order of $10^{-9}$. Therefore, the entanglement entropy in this phase seems to be finite. However, it is difficult to verify this property since the model parameters in this phase lie in a finite interval. Also, the phase is surrounded by critical lines that drive the entanglement entropy to high values.
@@
Moving to the $\mathcal{PT}$-unbroken topological phase, we note that the correlation spectrum is composed by eigenvalues close to $1$ and 0, with some of them being slightly bigger than $1$ or smaller then 0, as well as a conjugate pair of the form $1/2\pm\text{i}\alpha$, see panel (b) in Fig. 14 where the parameters are set to $W=1.7,V=0.5$. As the lattice size increases, we note that the eigenvalues tend to $1$ or 0, while $\alpha$ decreases slowly and seems to stabilize with the length of the chain. The associated entanglement entropy is real and finite. Going deep into the phase by increasing $W$, as before, $\alpha$ decreases and the entanglement tends to $S_{N/2}\rightarrow 2\log{2}$, see both panels (a) and (b) in Fig. 15.
@@
In the $\mathcal{PT}$-unbroken trivial phase, the correlation spectrum is also composed mostly of 0’s and $1$’s, some of them slightly bigger than $1$ or smaller than 0. Also, some eigenvalues carry an imaginary part in the form of conjugate pairs, see panel (d) in Fig. 14 for the model parameters $W=0.5$ and $V=1.8$. Interestingly, similarly to previous cases, both real and imaginary part of the eigenvalues vary slowly with the length of the chain. Nevertheless, deep inside this phase (large $V$), the eigenvalues tend to 0 or $1$ and produces zero entanglement, as expected from a trivial phase; see panel (b) in Fig. 15.
@@
Let us now discuss the exceptional phase. In this phase we observe an intricate behavior of the correlation spectrum, characterized by various complex eigenvalues, see panel (c) in Fig. 14 where the parameters are set to V=0.5,W=1.2V=0.5,\,W=1.2. This behavior is reflected in an intricate evolution of the entanglement entropy, see panels (a) and (b) in Fig. 15. Interestingly, the entanglement entropy detects the presence of the real exceptional points (indicated by dashed lines in Fig. 15). For clarity, we only pick a few values of $W$ between the exceptional points. Note that in the thermodynamic limit the entanglement entropy is ill defined in the exceptional phase, since it is filled with exceptional points. For finite systems, we expect a scaling of $S_{N/2}$. This is numerically challenging and goes beyond the scope of this paper.